A sample of nitrogen gas exerts a pressure of 9.80 atm at 32 °C. What would its temperature be (in °C) when its pressure is increased to 11.2 atm?
1
Expert's answer
2013-02-28T10:31:22-0500
Modified Clapeyron-Mendeleev law equation can be used for answering this question:
PV=nRT
P1V1=nRT1
P2V2=nRT2

P1/T1=P2/T2
T2=P2/(P1/T1)

So the answer is T2=11.2 atm*32 °C /9.80 atm = 36,6 C